California​ Gardens, Inc.,​ prewashes, shreds, and distributes a variety of salad mixes in​ 2-pound bags. Doug​ Voss, Operations​ VP, is considering a new​ Hi-Speed shredder to replace the old​ machine, referred to in the shop as​ "Clunker." Hi-Speed will have a fixed cost of

​$80,000

per month and a variable cost of

​$1.30

per bag. Clunker has a fixed cost of only

​$36,000

per​ month, but a variable cost of

​$1.80.

Selling price is

​$2.60

per bag.

​a) What is the crossover point in units​ (point of​ indifference) for the​ processes?

The crossover point is

nothing

units. ​(Round your response to the nearest whole​ number.)

​b) What is the monthly profit or loss if the company changes to the​ Hi-Speed shredder and sells

70,000

bags per​ month?

The monthly profit or loss if the company changes to the​ Hi-Speed shredder and sells

70,000

bags per month is

​$nothing.

​(Round your response to the nearest dollar and include a minus sign if​ necessary.)

​c) What is the monthly profit or loss if the company stays with Clunker and sells

70,000

bags per​ month?

The monthly profit or loss if the company stays with Clunker and sells

70,000

bags per month is

​$nothing.

​(Round your response to the nearest dollar and include a minus sign if​ necessary.)
Answer #1

1. INDIFFERENCE POINT = (FIXED COST A - FIXED COST B) / (VARIABLE COST B - VARIABLE COST A)


FC A = 80000
VC A = 1.3
FC B = 36000
VC B = 2.6

INDIFFERENCE POINT(A TO B) = (80000 - 36000) / (2.6 - 1.3) = 33846

2.FIXED COST = 80000
VARIABLE COST = 1.3
PRICE PER UNIT = 2.6
EXPECTED VOLUME = 70000

PROFIT = UNITS * CONTRIBUTION MARGIN - FIXED COST
PROFIT = (70000 * (2.6 - 1.3)) - 80000 = 11000

3. FIXED COST = 36000
VARIABLE COST = 1.8
PRICE PER UNIT = 2.6
EXPECTED VOLUME = 70000

PROFIT = UNITS * CONTRIBUTION MARGIN - FIXED COST
PROFIT = (70000 * (2.6 - 1.8)) - 36000 = 20000

** Leaving a thumbs-up would really help me out. Let me know if you face any problems.